OpenStudy (anonymous):

I got this problem wrong on my test. Can someone please walk me through it to show me what I did wrong? Determine two pairs of polar coordinates for the point (5, 5) with 0° ≤ θ < 360°.

OpenStudy (anonymous):

These are the steps. If you have questions asked me. My friend posted this question so..... x² + y² = r² (5)² + (-5)² = r² 25 + 25 = r² 50 = r² √50 = r 5√2 = r. To solve Θ: y/x = sin Θ / cos Θ y/x = tan Θ -5/5 = tan Θ -1 = tan Θ Θ = arc tan (-1). Θ = arc tan (-1) = 315°, and Θ = arc tan (-1) = (315° - 180°) = 135°. Answer : (5√2, 315°), and (5√2, 135°).

OpenStudy (anonymous):

btw I think you wrote your coordinates wrong.

OpenStudy (anonymous):

I got the exact same thing as this except it was not the correct answer

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

http://prntscr.com/41c9ii

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

are you going with byu?

OpenStudy (anonymous):

I had already answered b I am just wondering what i did wrong because my final is next week

OpenStudy (anonymous):

BYU?

OpenStudy (anonymous):

what is byu

OpenStudy (anonymous):

never mind. It's an online course university. I thought you were going with that university...because the question format looked similar.

OpenStudy (anonymous):

No I am not sure what it is.. I'm not in university I am in highschool

OpenStudy (anonymous):

no I'm in high school too but I'm taking an online course with a university ...

OpenStudy (anonymous):

except the coordinate on that one is negative

OpenStudy (anonymous):

only difference is the points are -5,5

OpenStudy (anonymous):

which is different but same sol.

OpenStudy (anonymous):

the thing I don't understand is -5^2 and 5^2 gets the same solution. 25

OpenStudy (anonymous):

Yeah it will get the same solution for the beginning part to find r because (-5)^2=5^2

OpenStudy (anonymous):

but it messed up the coordinates for theta

OpenStudy (anonymous):

I still think the answer is Answer : (5√2, 315°), and (5√2, 135°).

OpenStudy (anonymous):

ur right give me a sec.

OpenStudy (anonymous):

@axhleem what do you think?